OpenStudy (howard-wolowitz):

CAn someone show me how to do this:

OpenStudy (campbell_st):

so here is an example data set 1 1, 2, 3, what is the mean...? data set 2 9, 10, 11 what is the mean...? stanadard deviation uses the gaps between the data... ( this a a very simple explanation).... do the gaps change....?

OpenStudy (howard-wolowitz):

how could they not change?

OpenStudy (campbell_st):

well the mean increases by 8 since every score was increased by 8... does that make sense...?

OpenStudy (howard-wolowitz):

yes it does

OpenStudy (campbell_st):

ok... now for the standard deviation.... data set 1: 1, 2, 3 mean 2 (1 - 2), (2 - 2), (3 - 2) data set 2 9, 10, 11 mean 10 (9-10), (10 - 10), (11 - 10) do the gaps between the mean and each data value change...

OpenStudy (howard-wolowitz):

yes they do

OpenStudy (campbell_st):

well actually they don't data set 1, using score - mean data values 1, 2, 3, (1 - 2)= -1 (2 - 2) = 0 and (3 - 2) = 1 data set 2 9, 10, 11 (9 - 10) = -1 (10 - 10) = 0 and ( 11 - 10) = 1 so do the gaps change...?

OpenStudy (howard-wolowitz):

no. but I thought they did

OpenStudy (campbell_st):

the standard deviation is a measure of spread... so if every score is increased by the same value, then the gaps between the data don't change... and in the other question, again standard deviation is a measure of spread the smaller the standard deviation, the closer the data is... the larger the standard deviation the more spread the data... hope it helps